OpenStudy (pokiedokie):

Write the equation in slope-intercept form. slope=1/2 and the y intercept is (0,6) Walk me through this please!!

OpenStudy (anonymous):

The slope-intercept form is: \( y=mx+b\), where \(m\) is the slope. So \(m=\frac{1}{2}\). We also have a point \((0,6)\). This point is your x-intercept and y-intercept. \(x=0\) and \(y=6\) So we have \(y\), \(x\), and \(m\), but we don't have the \(b\). Plug in what's given to the slope-intercept form to solve for \(b\). \( y=mx+b\) \(6=(\frac{1}{2})(0)+b\) \(b=6\).

OpenStudy (pokiedokie):

y = 1/2 + 6?? or do I incorporate the 0

OpenStudy (anonymous):

We already have the things we need, so we should now write it in \({slope-intercept form\). \(\huge m=\frac {1}{2}\) and \(\huge b=6\). \(\huge y=mx+b\) \(\huge y=\frac {1}{2}x+6\)

OpenStudy (pokiedokie):

ahhh okay I see, that's what I thought I was just unsure if I add in the 0 or not. Can you please help me on another equation? I'm having trouble on the placements